### v1

@@ 3.1.2 Techniques related to safety and their relations to other aspects of the AI Act
Stricter forward alignment: Prompt injection happens when a user deliberately inputs text to modify the behavior of a Large Language Model (LLM), aiming to circumvent safeguards for producing misinformation, offensive content, or extracting private information [54]. To counteract these attacks, model developers are implementing stronger built-in protections by enhancing forward alignment through training on selected, pre-aligned datasets and further refining the models via Reinforcement Learning informed by Human Feedback (RLHF) [107][105]. This method of training, which indirectly incorporates human judgment to fine-tune models, helps ensure that LLMs are more aligned with human values and less prone to undesirable outputs. OpenAI’s GPT-4, for instance, underwent fine-tuning with RLHF, resulting in a reduced propensity for generating harmful or inaccurate content [41].
@@
Prompt instruction for LLM: LLM instructions can guide the model in handling user inputs with caution. By adding detailed instructions to the input prompt, models can be prepped to recognize and respond appropriately to content that might lead to a jailbreak, where the model’s behavior deviates from intended safeguards [114]. Strategically placing user input before these instructions can leverage the model’s tendency to prioritize recent instructions, enhancing adherence to safety protocols [108]. Furthermore, surrounding the instructions with random characters or specific HTML tags can signal to the model the distinction between system directives and user-provided prompts. This method helps delineate the boundaries between what the model perceives as commands to follow versus input to analyze, thereby improving the model’s ability to navigate and process inputs more securely and effectively. Well-designed prompt instructions that prioritize fairness, sensitivity to context, and inclusive language can contribute to increased fairness in LLMs. However, it’s essential to carefully consider the design and implementation of prompt instructions to avoid unintended consequences that could decrease fairness.
@@
Stricter backward alignment: Model providers are increasingly developing and integrating robust mechanisms by focusing on enhanced backward alignment. This involves rigorous training and evaluation processes using benchmark datasets specifically designed for this purpose or through filters that carefully monitor both the inputs received and outputs generated by a secured LLM. A notable strategy involves assessing the responses of LLMs to various prompts, which can help in identifying prompts that are potentially adversarial in nature, thereby preventing misuse of the model [105].
@@
Interpretability-based solutions: Interpretability-based solutions targeting attacks on LLMs involve using outlier detection techniques to monitor the models’ prediction trajectories. This approach is grounded in the observation that how a model processes and predicts outcomes based on inputs can reveal a lot about the nature of those inputs. Specifically, researchers have found that by closely examining the “prediction trajectory” - the series of predictions an LLM makes as it processes input-it’s possible to identify inputs that are anomalous or potentially malicious [107].
The concept of a “tuned lens” refers to a method or tool designed to scrutinize the model’s prediction path. When this lens is applied to analyze how an LLM responds to various inputs, it can highlight when the model’s predictions deviate significantly from what is expected under normal circumstances. Anomalous inputs, such as those designed to manipulate the model or trigger undesired behavior, often lead to unusual prediction trajectories. By detecting these outliers, interpretability-based solutions can effectively flag potentially harmful inputs before they impact the model’s output, enhancing the security and reliability of LLMs in handling a wide range of data [115].
@@ 3.2.1 Overview
AI transparency is crucial for upholding the integrity and comprehensibility of AI systems [116, 117]. It involves straightforward communication about the functions, decision-making processes, and constraints of these systems. This level of openness enables users to understand how decisions are made and help build trust. Furthermore, AI transparency includes explainability, which requires providing clear explanations of how AI processes inputs to produce outputs. This aspect is essential for enhancing user confidence and ensuring that AI operations comply with regulations designed to protect rights and prevent biases. In this subsection, we explore transparency in AI by discussing the roles of explainability and information disclosure.
@@
-Towards transparency through explainability: AI explainability is crucial for understanding how AI models make decisions, ensuring transparency, accountability, and trust in AI systems. Several techniques can help achieve AI transparency by providing clear, understandable explanations for AI model behavior. Regarding model interpretability methods, a model that is interpretable, such as a decision tree, a linear regression, or a rule-based system, is transparent, and the model itself is intrinsically understandable [118]. Techniques such as the SHapley Additive exPlanations (SHAP) and Local Interpretable Model-agnostic Explanations (LIME) offer insights into which features are most important for the model’s predictions [119]. These feature importance techniques can help trace back the decision-making process by highlighting the contribution of each input feature to the final decision. Some works utilized methods, such as counterfactual explanations [120] and prototypes [121], providing insights by presenting specific examples the model uses for making decisions. Specifically, counterfactual explanations show how slight changes to the input can lead to different predictions, helping trace the decision boundaries of the model. Prototypes, on the other hand, identify representative examples from the dataset that are influential in the model’s learning. In addition, visualizations can offer intuitive insights into complex models, especially deep learning models [122]. Techniques like saliency maps, which highlight the parts of the input (e.g., pixels in an image) that are most influential for a prediction, can help trace the model’s focus and reasoning process [123]. Techniques that decompose model predictions into understandable components. For instance, Layer-wise Relevance Propagation (LRP) decomposes the output decision back to the input features to identify the relevance of each feature in the decision-making process [124]. Lastly, tools and platforms allow users to interact with AI models, query them with different inputs and observe the outputs [125]. This hands-on approach can help users understand the model’s behavior in different scenarios and trace its decision-making process more practically and intuitively.
@@
-Explainability for Large Language Models: The techniques for enhancing Large Language Models (LLMs) explainability are categorized into two main paradigms: traditional fine-tuning methods and prompting-based methods [126]. Traditional fine-tuning methods encompass various strategies to enhance the interpretability and understanding of LLMs. Feature attribution methods form one facet of this approach, discerning the significance of different parts of input data, such as words or tokens within a sentence, for the model’s predictions. Techniques like Integrated Gradients [127] utilize gradients to gauge the importance of each input feature by tracing their impact on the model’s output. Surrogate models offer another avenue, employing simpler and more interpretable models to approximate the intricate behavior of complex LLMs [128]. While they provide insights into decision-making processes, they may not capture the full complexity of the original model. Representation analysis delves into understanding the learned representations of the model, often through visualization or clustering of embeddings to identify patterns or similarities captured by the model.
@@
Prompting-based methods leverage the generative capabilities of LLMs for explanation generation [126]. By prompting the model to articulate reasons for its decisions, this method sheds light on its underlying reasoning. Counterfactual explanations involve altering inputs slightly to observe changes in the model’s predictions, elucidating crucial aspects of the decision-making process [129]. Prompt engineering involves crafting prompts to elicit more interpretable or explainable responses from the model. These prompts may encourage the model to articulate its reasoning process or highlight key factors influencing its decisions, contributing to a deeper understanding of the model’s behavior [130].
@@
-Information disclosure: Advances in generative AI have led to AI-generated media that is nearly indistinguishable from human-created content. This has underscored the need for standardized accountability reports as a valuable measure to improve the transparency of AI systems across the industry [131]. According to the GDPR [3], providers of AI systems interacting with humans must ensure that people are informed of their interaction with AI unless it’s obvious from the context. Moreover, AI systems generating or manipulating content resembling real people, objects, or events (i.e., deepfakes) should explicitly disclose that the content is artificially created or manipulated.
@@ 3.2.2 Transparency versus safety, Environmental sustainability, non-discrimination, and traceability
While transparency strengthens traceability and supports non-discrimination by uncovering biases, it can conflict with safety and privacy concerns. Additionally, it promotes environmental sustainability but may reveal operational details. Thus, the ensuing discussion will elucidate the delicate balance required between transparency and the other principles to navigate the trade-offs and harmonize ethical AI deployment. As summarized in Table 4, the methods to augment transparency are categorized [132] into five categories: Dimension Reduction and Visualization, Feature-based Techniques, Example-based Techniques, Distributed Learning, and Pre-existing Model-based Methods. This subsection explores and discusses in more detail the approaches that can influence AI transparency under the six categories shown in Table 4.
@@
[133] [134] Dimension reduction & visualization PCA, LDA Visualizes feature distributions and their influence on model decisions, helping to understand model behavior and detect potential biases. Traceability(+), non-discrimination(+), environmental- sustainability(+)
[135] [136] [137] t-SNE, UMAP What-If
@@
[138] [139] [140] [141] Feature based methods ICE, LIME, SHAP, InterpretML Explains prediction outcomes by analyzing feature importance and their individual contributions to model decisions. Traceability(+) non-discrimination(+)
@@
[132] [142] [143] [144] [145] [137] Example based methods Anchors, CEM, KT SHAP, Dice, Alibi, What-If Demonstrates how specific changes in input conditions affect model predictions, making decision processes more interpretable. Traceability(+) non-discrimination(+)
@@
[146] [147] Distributed learning methods Federated Learning, Swam Learning Enables transparent model behavior adjustment at each network node while maintaining data privacy through limited sharing. Safety(-+), non-discrimination(+), environmental- sustainability(-)
@@
[100] Pre-existing model based methods Machine Unlearning Updates model behavior by modifying features or removing specific data influences. Non-discrimination(+), environmental sustainability(+)
@@
Dimension reduction and visualization: These techniques are essential for achieving data understanding and interpretability, thereby promoting transparent AI models. They simplify the complexity of large datasets, providing clearer insights and more transparent data narratives that can be integrated into AI algorithms for decision-making. For instance, Principal Component Analysis (PCA) [148] and Linear Discriminant Analysis (LDA) [134] can reduce high-dimensional features to low-dimensional spaces for feature space visualization, roughly explaining a model’s decisions. Although these transformations may obscure the original features’ meanings, the importance of features in model decisions can still be prioritized based on their similarity to PCA loadings [148]. Visualization techniques like t-SNE and UMAP offer a visual understanding of the model’s decision boundaries and the distribution of the training data [135, 136, 137].
@@
These techniques also positively impact the development of non-discriminative and environmentally sustainable AI with enhanced traceability. Specifically, PCA and LDA can unintentionally amplify biases if the principal components or discriminants used for dimensionality reduction capture biases inherent in the data [148, 134]. In contrast, t-SNE [135] and UMAP [136] preserve local structures and neighborhood relationships, helping to reveal and mitigate biases to ensure equitable treatment of all individuals and groups. Regarding environmental sustainability, dimensionality reduction with PCA and LDA can make model training more computationally efficient, leading to lower energy consumption and a smaller carbon footprint. Visualization techniques like t-SNE and UMAP can also reduce computational costs by providing insights that lead to more efficient model training and fewer iterations. Additionally, t-SNE and UMAP maintain relationships between data points in a lower-dimensional space, facilitating the traceability of how groups of data points are related to each other and potential biases, thus enhancing the traceability of decisions within the AI system.
@@
Feature-based methods: Deployed during the Deployment and Maintenance stage, these techniques elucidate the contribution of individual features to the outcomes of AI models. They allow for the dissection of model decisions, facilitating an understanding of how inputs are transformed into outputs. Feature-based methods such as ICE (i.e., Individual Conditional Expectation) [138], LIME (i.e., Local Interpretable Model-Agnostic Explanations) [139], SHAP (i.e., Shapley Additive explanations) [140], and Interpret-ML [141] significantly influence factors like non-discrimination and traceability in the realm of AI transparency.
@@
Example-based methods: Throughout the deployment and maintenance stage, these techniques focus on providing explanations for individual predictions made by AI systems. They support transparency by enabling stakeholders to query and receive understandable explanations for specific AI decisions. Example-based methods such as Anchors [132], Contrastive Explanation Method (CEM) [142], Kernel Tree SHAP [143], Diverse counterfactual explanations (Dice) [144], and Alibi [145] positively impact non-discrimination and traceability in the context of AI system transparency:
@@
Distributed learning methods: Distributed learning methods offer decentralized AI solutions that align with inherently decentralized data structures and comply with data privacy and security regulations. These methods, like federated learning [147] and swarm learning [146], ensure that raw data does not need to be exchanged, maintaining proper transparency for all participants. Specifically, swarm learning provides secure, transparent, and fair onboarding for decentralized network members without requiring a central custodian [146].
While improving AI system transparency, distributed learning methods have mixed impacts: they positively affect non-discrimination, negatively affect environmental sustainability, and have complex effects on safety. Regarding non-discrimination, collaborative fairness can be achieved by adjusting the performance of models allocated to each participant based on their contributions [149]. However, distributed learning methods are not environmentally friendly in AI development. Qiu et al. [150] found that training time for distributed learning is significantly longer than for centralized training due to the lower computational capabilities of participants’ devices. The impact on safety is multifaceted. A survey highlights the positive impact on data privacy through distributed learning in AI development but also points out vulnerabilities to various attacks, such as poisoning, backdoor, Generative Adversarial Network (GAN)-based attacks, and inference-based attacks [151].
Pre-existing model-based methods: Unlike the aforementioned post-hoc explainable AI methods, this category encompasses approaches that ensure AI model transparency through modifications or adjustments to the model’s capabilities. Specifically, machine unlearning [100] ensures the user’s “right to be forgotten” by enabling the model to erase the knowledge associated with a specific user. As discussed in section 3.1.2, this method also has a positive effect on non-discrimination but positively impacts the environmental cost of AI.

### v2

@@ 3.1.2 Techniques related to safety and their relations to other aspects of the AI Act
Stricter forward alignment: Prompt injection happens when a user deliberately inputs text to modify the behavior of a Large Language Model (LLM), aiming to circumvent safeguards for producing misinformation, offensive content, or extracting private information (?). To counteract these attacks, model developers are implementing stronger built-in protections by enhancing forward alignment through training on selected, pre-aligned datasets and further refining the models via Reinforcement Learning informed by Human Feedback (RLHF) (?)(?). This method of training, which indirectly incorporates human judgment to fine-tune models, helps ensure that LLMs are more aligned with human values and less prone to undesirable outputs. OpenAI’s GPT-4, for instance, underwent fine-tuning with RLHF, resulting in a reduced propensity for generating harmful or inaccurate content (?).
@@
Prompt instruction for LLM: LLM instructions can guide the model in handling user inputs with caution. By adding detailed instructions to the input prompt, models can be prepped to recognize and respond appropriately to content that might lead to a jailbreak, where the model’s behavior deviates from intended safeguards (?). Strategically placing user input before these instructions can leverage the model’s tendency to prioritize recent instructions, enhancing adherence to safety protocols (?). Furthermore, surrounding the instructions with random characters or specific HTML tags can signal to the model the distinction between system directives and user-provided prompts. This method helps delineate the boundaries between what the model perceives as commands to follow versus input to analyze, thereby improving the model’s ability to navigate and process inputs more securely and effectively. Well-designed prompt instructions that prioritize fairness, sensitivity to context, and inclusive language can contribute to increased fairness in LLMs. However, it’s essential to carefully consider the design and implementation of prompt instructions to avoid unintended consequences that could decrease fairness.
@@
Stricter backward alignment: Model providers are increasingly developing and integrating robust mechanisms by focusing on enhanced backward alignment. This involves rigorous training and evaluation processes using benchmark datasets specifically designed for this purpose or through filters that carefully monitor both the inputs received and outputs generated by a secured LLM. A notable strategy involves assessing the responses of LLMs to various prompts, which can help in identifying prompts that are potentially adversarial in nature, thereby preventing misuse of the model (?).
@@
Interpretability-based solutions: Interpretability-based solutions targeting attacks on LLMs involve using outlier detection techniques to monitor the models’ prediction trajectories. This approach is grounded in the observation that how a model processes and predicts outcomes based on inputs can reveal a lot about the nature of those inputs. Specifically, researchers have found that by closely examining the “prediction trajectory” - the series of predictions an LLM makes as it processes input-it’s possible to identify inputs that are anomalous or potentially malicious (?).
The concept of a “tuned lens” refers to a method or tool designed to scrutinize the model’s prediction path. When this lens is applied to analyze how an LLM responds to various inputs, it can highlight when the model’s predictions deviate significantly from what is expected under normal circumstances. Anomalous inputs, such as those designed to manipulate the model or trigger undesired behavior, often lead to unusual prediction trajectories. By detecting these outliers, interpretability-based solutions can effectively flag potentially harmful inputs before they impact the model’s output, enhancing the security and reliability of LLMs in handling a wide range of data (?).
@@ 3.2.1 Overview
AI transparency is crucial for upholding the integrity and comprehensibility of AI systems (?, ?). It involves straightforward communication about the functions, decision-making processes, and constraints of these systems. This level of openness enables users to understand how decisions are made and help build trust. Furthermore, AI transparency includes explainability, which requires providing clear explanations of how AI processes inputs to produce outputs. This aspect is essential for enhancing user confidence and ensuring that AI operations comply with regulations designed to protect rights and prevent biases. In this subsection, we explore transparency in AI by discussing the roles of explainability and information disclosure.
@@
-Towards transparency through explainability: AI explainability is crucial for understanding how AI models make decisions, ensuring transparency, accountability, and trust in AI systems. Several techniques can help achieve AI transparency by providing clear, understandable explanations for AI model behavior. Regarding model interpretability methods, a model that is interpretable, such as a decision tree, a linear regression, or a rule-based system, is transparent, and the model itself is intrinsically understandable (?). Techniques such as the SHapley Additive exPlanations (SHAP) and Local Interpretable Model-agnostic Explanations (LIME) offer insights into which features are most important for the model’s predictions (?). These feature importance techniques can help trace back the decision-making process by highlighting the contribution of each input feature to the final decision. Some works utilized methods, such as counterfactual explanations (?) and prototypes (?), providing insights by presenting specific examples the model uses for making decisions. Specifically, counterfactual explanations show how slight changes to the input can lead to different predictions, helping trace the decision boundaries of the model. Prototypes, on the other hand, identify representative examples from the dataset that are influential in the model’s learning. In addition, visualizations can offer intuitive insights into complex models, especially deep learning models (?). Techniques like saliency maps, which highlight the parts of the input (e.g., pixels in an image) that are most influential for a prediction, can help trace the model’s focus and reasoning process (?). Techniques that decompose model predictions into understandable components. For instance, Layer-wise Relevance Propagation (LRP) decomposes the output decision back to the input features to identify the relevance of each feature in the decision-making process (?). Lastly, tools and platforms allow users to interact with AI models, query them with different inputs and observe the outputs (?). This hands-on approach can help users understand the model’s behavior in different scenarios and trace its decision-making process more practically and intuitively.
@@
-Explainability for Large Language Models: The techniques for enhancing Large Language Models (LLMs) explainability are categorized into two main paradigms: traditional fine-tuning methods and prompting-based methods (?). Traditional fine-tuning methods encompass various strategies to enhance the interpretability and understanding of LLMs. Feature attribution methods form one facet of this approach, discerning the significance of different parts of input data, such as words or tokens within a sentence, for the model’s predictions. Techniques like Integrated Gradients (?) utilize gradients to gauge the importance of each input feature by tracing their impact on the model’s output. Surrogate models offer another avenue, employing simpler and more interpretable models to approximate the intricate behavior of complex LLMs (?). While they provide insights into decision-making processes, they may not capture the full complexity of the original model. Representation analysis delves into understanding the learned representations of the model, often through visualization or clustering of embeddings to identify patterns or similarities captured by the model.
@@
Prompting-based methods leverage the generative capabilities of LLMs for explanation generation (?). By prompting the model to articulate reasons for its decisions, this method sheds light on its underlying reasoning. Counterfactual explanations involve altering inputs slightly to observe changes in the model’s predictions, elucidating crucial aspects of the decision-making process (?). Prompt engineering involves crafting prompts to elicit more interpretable or explainable responses from the model. These prompts may encourage the model to articulate its reasoning process or highlight key factors influencing its decisions, contributing to a deeper understanding of the model’s behavior (?).
@@
-Information disclosure: Advances in generative AI have led to AI-generated media that is nearly indistinguishable from human-created content. This has underscored the need for standardized accountability reports as a valuable measure to improve the transparency of AI systems across the industry (?). According to the GDPR (?), providers of AI systems interacting with humans must ensure that people are informed of their interaction with AI unless it’s obvious from the context. Moreover, AI systems generating or manipulating content resembling real people, objects, or events (i.e., deepfakes) should explicitly disclose that the content is artificially created or manipulated.
@@ 3.2.2 Transparency versus safety, Environmental sustainability, non-discrimination, and traceability
While transparency strengthens traceability and supports non-discrimination by uncovering biases, it can conflict with safety and privacy concerns. Additionally, it promotes environmental sustainability but may reveal operational details. Thus, the ensuing discussion will elucidate the delicate balance required between transparency and the other principles to navigate the trade-offs and harmonize ethical AI deployment. As summarized in Table 4, the methods to augment transparency are categorized (?) into five categories: Dimension Reduction and Visualization, Feature-based Techniques, Example-based Techniques, Distributed Learning, and Pre-existing Model-based Methods. This subsection explores and discusses in more detail the approaches that can influence AI transparency under the six categories shown in Table 4.
@@
(?, ?) Dimension reduction & visualization PCA, LDA Visualizes feature distributions and their influence on model decisions, helping to understand model behavior and detect potential biases. Traceability (+), non-discrimination (+), environmental sustainability (+).
(?, ?, ?) t-SNE, UMAP What-If
@@
(?, ?, ?, ?) Feature based methods ICE, LIME, SHAP, InterpretML Explains prediction outcomes by analyzing feature importance and their individual contributions to model decisions. Traceability (+), non-discrimination (+).
@@
(?, ?, ?, ?, ?, ?) Example based methods Anchors, CEM, KT SHAP, Dice, Alibi, What-If Demonstrates how specific changes in input conditions affect model predictions, making decision processes more interpretable. Traceability (+), non-discrimination (+).
@@
(?, ?) Distributed learning methods Federated Learning, Swam Learning Enables transparent model behavior adjustment at each network node while maintaining data privacy through limited sharing. Safety (-+), non-discrimination (+), environmental sustainability (-).
@@
(?) Pre-existing model based methods Machine Unlearning Updates model behavior by modifying features or removing specific data influences. Non-discrimination (+), environmental sustainability (+).
@@
Dimension reduction and visualization: These techniques are essential for achieving data understanding and interpretability, thereby promoting transparent AI models. They simplify the complexity of large datasets, providing clearer insights and more transparent data narratives that can be integrated into AI algorithms for decision-making. For instance, Principal Component Analysis (PCA) (?) and Linear Discriminant Analysis (LDA) (?) can reduce high-dimensional features to low-dimensional spaces for feature space visualization, roughly explaining a model’s decisions. Although these transformations may obscure the original features’ meanings, the importance of features in model decisions can still be prioritized based on their similarity to PCA loadings (?). Visualization techniques like t-SNE and UMAP offer a visual understanding of the model’s decision boundaries and the distribution of the training data (?, ?, ?).
@@
These techniques also positively impact the development of non-discriminative and environmentally sustainable AI with enhanced traceability. Specifically, PCA and LDA can unintentionally amplify biases if the principal components or discriminants used for dimensionality reduction capture biases inherent in the data (?, ?). In contrast, t-SNE (?) and UMAP (?) preserve local structures and neighborhood relationships, helping to reveal and mitigate biases to ensure equitable treatment of all individuals and groups. Regarding environmental sustainability, dimensionality reduction with PCA and LDA can make model training more computationally efficient, leading to lower energy consumption and a smaller carbon footprint. Visualization techniques like t-SNE and UMAP can also reduce computational costs by providing insights that lead to more efficient model training and fewer iterations. Additionally, t-SNE and UMAP maintain relationships between data points in a lower-dimensional space, facilitating the traceability of how groups of data points are related to each other and potential biases, thus enhancing the traceability of decisions within the AI system.
@@
Feature-based methods: Deployed during the Deployment and Maintenance stage, these techniques elucidate the contribution of individual features to the outcomes of AI models. They allow for the dissection of model decisions, facilitating an understanding of how inputs are transformed into outputs. Feature-based methods such as ICE (i.e., Individual Conditional Expectation) (?), LIME (i.e., Local Interpretable Model-Agnostic Explanations) (?), SHAP (i.e., Shapley Additive explanations) (?), and Interpret-ML (?) significantly influence factors like non-discrimination and traceability in the realm of AI transparency.
@@
Example-based methods: Throughout the deployment and maintenance stage, these techniques focus on providing explanations for individual predictions made by AI systems. They support transparency by enabling stakeholders to query and receive understandable explanations for specific AI decisions. Example-based methods such as Anchors (?), Contrastive Explanation Method (CEM) (?), Kernel Tree SHAP (?), Diverse counterfactual explanations (Dice) (?), and Alibi (?) positively impact non-discrimination and traceability in the context of AI system transparency:
@@
Distributed learning methods: Distributed learning methods offer decentralized AI solutions that align with inherently decentralized data structures and comply with data privacy and security regulations. These methods, like federated learning (?) and swarm learning (?), ensure that raw data does not need to be exchanged, maintaining proper transparency for all participants. Specifically, swarm learning provides secure, transparent, and fair onboarding for decentralized network members without requiring a central custodian (?).
While improving AI system transparency, distributed learning methods have mixed impacts: they positively affect non-discrimination, negatively affect environmental sustainability, and have complex effects on safety. Regarding non-discrimination, collaborative fairness can be achieved by adjusting the performance of models allocated to each participant based on their contributions (?). However, distributed learning methods are not environmentally friendly in AI development. Qiu et al. (?) found that training time for distributed learning is significantly longer than for centralized training due to the lower computational capabilities of participants’ devices. The impact on safety is multifaceted. A survey highlights the positive impact on data privacy through distributed learning in AI development but also points out vulnerabilities to various attacks, such as poisoning, backdoor, Generative Adversarial Network (GAN)-based attacks, and inference-based attacks (?).
Pre-existing model-based methods: Unlike the aforementioned post-hoc explainable AI methods, this category encompasses approaches that ensure AI model transparency through modifications or adjustments to the model’s capabilities. Specifically, machine unlearning (?) ensures the user’s “right to be forgotten” by enabling the model to erase the knowledge associated with a specific user. As discussed in section 3.1.2, this method also has a positive effect on non-discrimination but positively impacts the environmental cost of AI.